OpenStudy (anonymous):

What is the value of the expression w – 4 when w = –17? im pretty sure its 13 but im not sure

OpenStudy (theedwardsfamily):

nerp. 13

OpenStudy (theedwardsfamily):

*not 13

OpenStudy (anonymous):

-13? i have like 1 more

OpenStudy (theedwardsfamily):

w= -17 so... -17-4 = ?

OpenStudy (anonymous):

so its -13? or do we add it?

OpenStudy (theedwardsfamily):

its not -13 or 13. -17- 4 = 17+ 4 =?

OpenStudy (theedwardsfamily):

you turn ur answer into a negative

OpenStudy (theedwardsfamily):

so wht would ur answer be? @ilovefood123XD

OpenStudy (anonymous):

so its -21?

OpenStudy (anonymous):

-21?

OpenStudy (anonymous):

@TheEdwardsFamily

OpenStudy (theedwardsfamily):

yes. srry i was taking so long, lag

OpenStudy (theedwardsfamily):

your answer is -21

OpenStudy (theedwardsfamily):

good job!

OpenStudy (anonymous):

its okay and thanks can you help me with one more i just posted @TheEdwardsFamily

OpenStudy (theedwardsfamily):

ok